Suppose that the number of customers that enter a bank in an hour is a Poisson random variable, and suppose that P(X =0) = 0.02.

i. What is the probability that at most one person enter the bank in one hour?

ii. Determine the mean and standard deviation of X


1
Expert's answer
2021-07-01T17:49:15-0400

PMF of a Poisson distribution is "p(x;\lambda) = \frac{e^{-\lambda}\lambda^{x}} {x!} for 
x = 0, 1, 2, \cdots"

"P(x=0)= \frac{e^{-\lambda}\lambda^{0}} {0!}"

"e^{-\lambda} =0.02"

"\lambda=-\ln0.02"

i. "P(\le1)"

"P(x\le1)=P(x=0)+P(x=1)"

"=0.02+\frac{e^{\ln{0.02}}(-\ln0.02)} {1!}"

"=0.02+0.02\times3.9120"

"=0.09824"

The excel function =POISSON.DIST(1,-LN(0.02),TRUE) can be used to obtain the answer.

ii. Mean and standard deviation of x

The mean and variance of a Poisson distribution is "\lambda"

Thus,

The mean is "-\ln(0.02)=3.912" and

Standard deviation is "\sqrt{-\ln(0.02)}=\sqrt{3.912}=1.978"